How fast is an object moving if it has a centripetal acceleration of 460.8 m/s

2

460.8 m/s2

 and turns around in a circle with a radius of 5 m

5 m


1
Expert's answer
2021-10-25T15:24:40-0400

The centripetal acceleration for uniform circular motion is given as follows:


"a = \dfrac{v^2}{R}"

where "v" is the speed of the object (constant), "R = 5m" is the radius of the circle. Expressing "v" and substituting "a = 460.8m\/s^2", obtain:


"v = \sqrt{aR} = \sqrt{460.8m/s^2\cdot 5m} = 48m/s"

Answer. 48 m/s.